OpenStudy (clamin):

PLEASE HELP!! What number must you add to complete the square? x^2+8x=15 a.32 b. 4 c. 16 d. 8

OpenStudy (xapproachesinfinity):

first subtract 15 from both sides x^2+8x-15=0

OpenStudy (xapproachesinfinity):

you can factor that!

OpenStudy (clamin):

is it d??

OpenStudy (xapproachesinfinity):

so any brain storming here

OpenStudy (xapproachesinfinity):

oh my bad sorry i did not read the question clearly

OpenStudy (xapproachesinfinity):

forgot my proposal

OpenStudy (xapproachesinfinity):

you need to complete the square

OpenStudy (xapproachesinfinity):

focus on left hand side x^2+8x what can you add to make that a perfect square

OpenStudy (xapproachesinfinity):

x^2+2(4x)+ ....

OpenStudy (xapproachesinfinity):

you did say the answer before

OpenStudy (mathmale):

Complete the square: x^2+8x=15 Take HALF of the coefficient of 8x. Your response?

OpenStudy (xapproachesinfinity):

sorry i confused you :) my bad

OpenStudy (clamin):

4

OpenStudy (mathstudent55):

Now square it.

OpenStudy (clamin):

4*4=16

OpenStudy (mathmale):

Good. Now write +16 on both sides of your equation. Share your work, please.

OpenStudy (mathmale):

Write the whole equation, including the +16 on the left side and another +16 on the right side.

OpenStudy (clamin):

x^2+8x+16=15+16

OpenStudy (mathmale):

Very nice. Note that your x^2 + 8x + 16 is a perfect square. Can you write it as the square of a binomial? ( ? + ? ) ^2

OpenStudy (clamin):

(x +4)^2 is this right??

OpenStudy (mathstudent55):

correct what is 15 + 16? that goes on the right side

OpenStudy (mathstudent55):

\(x^2+8x+\color{red}{16}=15+16\) What did you add to complete the square?

OpenStudy (clamin):

x^2+ 8x +16 =31

OpenStudy (clamin):

@mathstudent55 @mathmale

OpenStudy (mathmale):

(x +4)^2 is this right?? yes, it is correct. Thus, you have (x+4)^2=31 Take the square root of both sides, being certain to include the plus-or-minus sign in front of your Sqrt(31)., Your result? Now solve this result for x. You MUST end up with two different roots.

OpenStudy (jiteshmeghwal9):

\[x^2+2.4.x=15\]here all u r missing is \(4^2\)therefore add \(4^2\) both the sides\[x^2+2.4.x+4^2=15+16\]\[(x+4)^2=31\]

OpenStudy (lhk416):

Take 8, divide it by 2, and square the result. The answer is 16. You need to add 16 to both sides to complete the square.